OpenStudy (anonymous):

I have the answer but I don't understand the process of subtracting the mixed number 2 1/10-1/4

OpenStudy (anonymous):

one way is to "borrow" another is to rewrite the first as an improper fraction \[2\tfrac{1}{10}=\frac{21}{10}\]

OpenStudy (anonymous):

that will give you \[\frac{21}{10}-\frac{1}{4}=\frac{42}{20}-\frac{5}{20}=\frac{37}{20}=1\tfrac{17}{20}\]\]

OpenStudy (anonymous):

Thank you, that was a great help.